sweetburger:

for fun question. Determine the partial derivative of cos(y+z) with respect to z using the definition of the derivative

ThisGirlPretty:

@Angle

ThisGirlPretty:

@Vocaloid

Vocaloid:

Intuitively I believe it would just be -sin(y+z) I just learned partials so I could be wrong ;;

Angle:

@vocaloid wouldn't you also have to multiply it by the derivative of the inside? (with respect to z)

Angle:

but to ask this with the definition of the derivative is a bit more difficult ^_^"

sweetburger:

Ok so yes your intuitive answer is correct @Vocaloid but the real fun comes with determining this using the definition of the derivative

sweetburger:

and @angle the derivative of the interior of cos(y+z) with respect to z is 1 so 1*-sin(y+z) is still just -sin(y+z)

sweetburger:

the goal is to use this \[\lim_{\Delta z \rightarrow 0}\frac{ f(x,y,z+\Delta z)-f(x,y,z) }{ \Delta z }\]

sillybilly123:

\(\dfrac{\partial [\cos (y + z)]}{\partial z} := \lim\limits_{\delta z \to 0}\dfrac{cos (y + z + \delta z) - \cos (y + z)}{\delta z}\) \(= \lim\limits_{\delta z \to 0}\dfrac{\cos (y + z) \cos \delta z - \sin (y + z) \sin \delta z - \cos (y + z)}{\delta z}\) Re-arrange: \(= \lim\limits_{\delta z \to 0} \left( \cos (y + z) \dfrac{\cos \delta z - 1}{\delta z} - \sin (y + z) \dfrac{\sin \delta z}{\delta z} \right)\) Limit of sum is sum of limits: \(= \cos (y + z) \lim\limits_{\delta z \to 0}\dfrac{\cos \delta z - 1}{\delta z} - \sin (y + z) \lim\limits_{\delta z \to 0} \dfrac{\sin \delta z}{\delta z} = \triangle\) Both these trig limits are well known: \( \left. \dfrac{\cos \delta z - 1}{\delta z} \right|_{\delta z \to 0} = 0\) \( \left. \dfrac{\sin \delta z }{\delta z} \right|_{\delta z \to 0} = 1\) So: \(\triangle = \cos (y + z) \times 0 - \sin (y + z) \times 1\) Link to the common trig limit + proofs: http://tutorial.math.lamar.edu/Classes/CalcI/ProofTrigDeriv.aspx